# Model Selection for Off-policy Evaluation: New Algorithms and Experimental Protocol

### v2

@@ Abstract
Holdout validation and hyperparameter tuning from data is a long-standing problem in offline reinforcement learning (RL). A standard framework is to use off-policy evaluation (OPE) methods to evaluate and select between different policies, but OPE methods either incur exponential variance (e.g., importance sampling) or have hyperparameters of their own (e.g., FQE and model-based). We focus on model selection for OPE itself, which is even more under-investigated. Concretely, we select among candidate value functions (“model-free”) or dynamics (“model-based”) to best assess the performance of a target policy. We develop: (1) new model-free and model-based selectors with theoretical guarantees, and (2) a new experimental protocol for empirically evaluating them. Compared to the model-free protocol in prior works, our new protocol allows for more stable generation and better control of candidate value functions in an optimization-free manner, and evaluation of model-free and model-based methods alike. We exemplify the protocol on Gym-Hopper, and find that our new model-free selector, LSTD-Tournament, demonstrates promising empirical performance.
@@ 1 Introduction
In this work, we make progress on the latter problem, namely model selection for OPE algorithms themselves. We lay out and study a dichotomy of two settings: in the model-based setting, evaluation algorithms build dynamics models to evaluate a target policy. Given the uncertainty of hyperparameters in model learning, we assume that multiple candidate models are given, and the task is to select one that best evaluates the performance of the target policy. In the model-free setting, evaluation algorithms only output value functions which we select from; see Figure 5 for a visualization of the pipeline. Notably, these selection algorithms should be hyperparameter-free themselves, to avoid further chicken-and-egg problems. Our contributions are:
@@
    New experiment protocol (Section 5): To empirically evaluate the selection algorithms, prior works often use FQE to prepare candidate Q-functions [61, 44], which suffer from unstable training [^1] For example, our preliminary investigation has found that FQE often diverges with CQL-trained policies [25], which is echoed by Nie et al. [44] in personal communications. and lack of control in the quality of the candidate functions. We propose a new experiment protocol, where the candidate value functions are induced from variations of the groundtruth environment; see Figure 5 for an illustration. This bypasses the caveats of FQE and allows for the computation of Q-values in an optimization-free and controllable manner. Moreover, the protocol can also be used to evaluate and compare estimators for the model-based setting. Implementation-wise, we use lazy evaluation and Monte-Carlo roll-outs to generate the needed Q-values. Combined with parallelization and caching, we reduce the computational cost and make the evaluation of new algorithms easier.
@@ 2 Preliminaries and Model Selection Problem
An MDP is specified by $(\mathcal{S},\mathcal{A},P,R,\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition dynamics, $R:\mathcal{S}\times\mathcal{A}\to[0,R_{\max}]$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $d_{0}$ is the initial state distribution. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ induces a distribution over random trajectories, generated as $s_{0}\sim d_{0}$, $a_{t}\sim\pi(\cdot|s_{t})$, $r_{t}=R(s_{t},a_{t})$, $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. We use $\operatorname{Pr}_{\pi}[\cdot]$ and $\mathbb{E}_{\pi}[\cdot]$ to denote such a distribution and the expectation thereof. The performance of a policy is defined as $J(\pi):=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, which is in the range of $[0,V_{\max}]$ where $V_{\max}:=R_{\max}/(1-\gamma)$.
@@
  • Model-based: Each OPE instance outputs an MDP $M_{i}$ and uses $J_{M_{i}}(\pi)$ as an estimate of $J(\pi)$. W.l.o.g. we assume $M_{i}$ only differs from $M^{\star}$ in the transition $P_{i}$, as handling different reward functions is straightforward. The task is to select $\hat{M}$ from :=$\mathcal{M}:=\{M_{i}\}_{i\in[m]}$, such that $J_{\hat{M}}(\pi)\approx J(\pi)$. Similar to the model-free case, we assume $M^{\star}\in\mathcal{M}$ in the derivations.
@@ 3.2 LSTD-Tournament
Besides the realizability of $Q^{\pi}$, the guarantee also depends on the invertibility of $A$, which can be viewed as a coverage condition, since $A$ changes with the data distribution $\mu$ [2, 3, 21]. In fact, in the on-policy setting, $\sigma_{\min}(A)$ can be shown to be lower-bounded away from 00 [40]; see Appendix C.5.
@@
The subscript “$i,j$” makes it clear that the feature is created based on $Q_{i}$ and $Q_{j}$ as candidates, and we will use similar conventions for all quantities induced from $\phi_{i,j}$, e.g., $A_{i,j},b_{i,j}$, etc. Obviously, $Q^{\pi}$ is linear in $\phi_{i,j}$ with $\theta^{\star}\in\{[1,0]^{\top},[0,1]^{\top}\}$. Therefore, to choose between $Q_{i}$ and $Q_{j}$, we can calculate the LSTDQ loss of $[1,0]^{\top}$ and $[0,1]^{\top}$ under feature $\phi_{i,j}$ and choose the one with smaller loss. For $\theta=[1,0]^{\top}$, we have $A_{i,j}\theta-b_{i,j}=$
@@ 4 New Model-Based Selectors
[…]
@@ 5.1 The Protocol
    Candidate model list =$\mathcal{M}=\{M_{i}\}_{i\in[m]}$.
@@
Our innovation here is to note that, for all the model-free algorithms we are interested in evaluating, they all access $Q_{M}^{\pi}$ exclusively through the value of $Q_{M}^{\pi}(s,a)$ and $Q_{M}^{\pi}(s^{\prime},\pi)$ for $(s,a,r,s^{\prime})$ in the offline dataset $\mathcal{D}$. That is, given $n$ data points in $\mathcal{D}$, we only need to know $2n$ scalar values about $Q_{M}^{\pi}$. Therefore, we propose to directly compute these values without explicitly representing $Q_{M}^{\pi}$, and each value can be easily estimated by averaging over multiple Monte-Carlo rollouts, i.e., $Q_{M}^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s,a_{0}=a]$ can be approximated by rolling out multiple trajectories starting from $(s,a)$ and taking actions according to $\pi$. For the model-based estimators proposed in Section 4, we need access to quantities in the form of $(\mathcal{T}_{M_{j}}^{\pi}Q_{M_{i}}^{\pi})(s,a)$ (see Appendix B). This value can also be obtained by Monte-Carlo simulation: (1) start in $(s,a)$ and simulate one step in $M_{j}$, then (2) switch to $M_{i}$, simulate from step 2 onwards and rollout the rest of the trajectory.
@@ 6 Exemplification of the Protocol
In this section we instantiate our protocol in the Gym Hopper environment to demonstrate its utility, while also providing preliminary empirical results for our algorithms.
@@ 6.1 Experiment Setup and Main Results
Our experiments will be based on the Hopper-v4 environment [8]. To create a variety of environments, we add different levels of stochastic noise in the transitions and change the gravity constant (see Appendix D.1). Each environment is then parameterized by the gravity constant $\mathbf{g}$ and noise level $\mathbf{n}$. We consider arrays of such environments as the set of candidate simulator $\mathcal{M}$: in most of our results, we consider a “gravity grid” (denoted using MF.G in the figures) $\mathcal{M}_{\mathbf{g}}:=\{M_{\mathbf{g}}^{0}\ldots,M_{\mathbf{g}}^{14}\}$ (fixed noise level, varying gravity constants from $-51$ to $-9$) and a “noise grid” (MF.N) $\mathcal{M}_{\mathbf{n}}:=\{M_{\mathbf{n}}^{0}\ldots,M_{\mathbf{n}}^{14}\}$ (fixed gravity constant, varying noise level from $10$ to $100$). Each array contains 15 environments, though some subsequent results may only involve a subset of them (Section 6.2). Some of these simulators will also be treated as groundtruth environment $M^{\star}$, which determines the groundtruth performance of target policies and produces the offline dataset $\mathcal{D}$.
@@ 6.3 Data Coverage
In the previous subsection, we have seen how multiple experiment units that only differ in $\mathcal{M}$ can provide useful insights. Here we show that we can also probe the methods’ sensitivity to data coverage by looking at experiment units that only differ in the dataset $\mathcal{D}$. In Figure 4R, we take a previous experiment setting ($\mathcal{M}_{\mathbf{g}}$) and isolate a particular target policy $\pi$; then, we create two datasets: (1) $\mathcal{D}_{\pi}$ sampled using $\pi$; (2) $\mathcal{D}_{\textrm{off}}$ sampled using a policy that is created to be very different from the target policies and offer very little coverage (see Appendix D.2). Then, we run the algorithm with $\lambda$ fraction of data from $\mathcal{D}_{\pi}$ combined with $(1-\lambda)$ from $\mathcal{D}_{\textrm{off}}$; as predicted by theory, most methods perform better with better coverage (large $\lambda$), and performance degrades as $\lambda$ goes to 00.
@@ Appendix A Other Related Works
Our theoretical guarantees state that the selected candidate function/model provides a $J(\pi)$ estimate that is close to the groundtruth, as if the algorithm knew which candidate is correct. This can be viewed as a form of adaptivity guarantees, which is the goal of several theoretical works on model selection. Su et al. [48] and Udagawa et al. [53] study adaptive model selection for OPE, but their approaches crucially rely on the importance sampling estimator, which we do not consider due to the focus on long-horizon tasks. Lee et al. [31] (who build on and improve upon the earlier works of Farahmand and Szepesvári [13] and Jiang et al. [22]) study model-selection of value functions, with a focus on the double-sampling issue and its relationship with Bellman completeness; these are also the very core consideration in our theoretical derivations. However, their approach treats the OPE instances (such as FQE) in a less black-box manner, and the goal is to select the “right” function class for the OPE algorithm (instead of directly selecting a final output function) from a nested series of classes, one of which is assumed to be Bellman-complete [6, 9] and have low generalization errors. [^4] Given the relationship between Bellman-completeness and bisimulation abstractions [9, Proposition 9], their setting is a natural generalization of Jiang et al. [22]’s setting of selecting a bisimulation from a nested series of state abstractions. This makes their approach less widely applicable than ours, and we cannot empirically evaluate their algorithm in our protocol as a consequence. Nevertheless, their results provide an interesting alternative and valuable insights to the model-selection problem, which are “complementary” to the BVFT line of work [59, 61] that we further develop. Similarly, Miyaguchi [39] considers the selection of Bellman operators, which can be instantiated as regression using different function classes, making their approach similar to Lee et al. [31] in spirit.
@@ Appendix B Details of Model-based Selectors
$\mathbb{E}_{(s,a,s^{\prime})\sim\mu,\tilde{s}\sim P(\cdot|s,a)}[d(s^{\prime},\tilde{s})],$ (12)
@@
Despite its wide use and simplicity [15, 43, 16], the method has major caveats: first, the distance metric $d(\cdot,\cdot)$ is a design choice. When the state is represented as a real-valued vector, it is natural to use the $\ell_{2}$ distance as $d(\cdot,\cdot)$, which changes if we simply normalize/rescale the coordinates. Second, the metric is biased for stochastic environments as discussed in prior works [19, 56], which we also demonstrate in the experiment section (Section 6); essentially this is a version of the double-sampling issue but for the model-based setting [5].

### v3

@@ Abstract
Holdout validation and hyperparameter tuning from data is a long-standing problem in offline reinforcement learning (RL). A standard framework is to use off-policy evaluation (OPE) methods to evaluate and select between different policies, but OPE methods either incur exponential variance (e.g., importance sampling) or have hyperparameters of their own (e.g., FQE and model-based). We focus on model selection for OPE itself, which is even more under-investigated. Concretely, we select among candidate value functions (“model-free”) or dynamics models (“model-based”) to best assess the performance of a target policy. We develop: (1) new model-free and model-based selectors with theoretical guarantees, and (2) a new experimental protocol for empirically evaluating them. Compared to the model-free protocol in prior works, our new protocol allows for more stable generation and better control of candidate value functions in an optimization-free manner, and evaluation of model-free and model-based methods alike. We exemplify the protocol on Gym-Hopper, and find that our new model-free selector, LSTD-Tournament, demonstrates promising empirical performance.
@@ 1 Introduction
In this work, we make progress on the latter problem, namely model selection for OPE algorithms themselves. We lay out and study a dichotomy of two settings: in the model-based setting, evaluation algorithms build dynamics models to evaluate a target policy. Given the uncertainty of hyperparameters in model learning, we assume that multiple candidate models are given, and the task is to select one that best evaluates the performance of the target policy. In the model-free setting, evaluation algorithms only output value functions which we select from; see Figure 1 (left) for a visualization of the pipeline. Notably, these selection algorithms should be hyperparameter-free themselves, to avoid further chicken-and-egg problems. Our contributions are:
@@
    New experiment protocol (Section 5): To empirically evaluate the selection algorithms, prior works often use FQE to prepare candidate Q-functions [64, 46], which suffer from unstable training [^1] For example, our preliminary investigation has found that FQE often diverges with CQL-trained policies [27], which is echoed by Nie et al. [46] in personal communications. and lack of control in the quality of the candidate functions. We propose a new experiment protocol, where the candidate value functions are induced from variations of the groundtruth environment; see Figure 1 (right) for an illustration. This bypasses the caveats of FQE and allows for the computation of Q-values in an optimization-free and controllable manner. Moreover, the protocol can also be used to evaluate and compare estimators for the model-based setting. Implementation-wise, we use lazy evaluation and Monte-Carlo roll-outs to generate the needed Q-values. Combined with parallelization and caching, we reduce the computational cost and make the evaluation of new algorithms easier.
@@
Pipeline in Practice (Model-free)
Experiment Protocol (Model-free)
Figure 1: An illustration of the pipeline in practice that motivates our research (left) and our proposed experimental protocol (right), both for the model-free setting; the pipeline for the model-based setting is analogous and not visualized. Left: Training algorithms are run on training data to generate candidate policies, and choosing among them is the primary model-selection (MS) task (see Section 1) which is not the focus of our work. These policies (e.g., $\pi$) become the target policies in OPE (e.g., $\pi$), since accurate OPE can help solve the primary MS problem. Different FQE instances (e.g., with different hyperparameters, such as neural architectures) are used to approximate $Q^{\pi}$, producing $\{Q_{i}\}$. The selector takes MS data and $\{Q_{i}\}$ as input and choose one of them to estimate $J(\pi)\equiv J_{M^{\star}}(\pi)$. Right: Illustration of our protocol for an experiment unit (Section 5). The target policies $\pi$ can be learned from separate training data but can also be produced in other ways, such as training on inaccurate models $M_{i}$, which can be realistic for practical scenarios with inaccurate simulators. $\{M_{i}\}$ is prepared by varying environment parameters. Monte-Carlo rollouts are used to generate the Q-values for the data points in the MS data, which avoids potentially unstable optimization in the OPE pipeline; this is the source of stability and controllability compared to the prior protocol that mimics the practical pipeline. For further discussion on the limitation of our protocol and the trade-offs, see Section 7.
@@ 2 Preliminaries and Model Selection Problem
An MDP is specified by $(\mathcal{S},\mathcal{A},P,R,\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition dynamics, $R:\mathcal{S}\times\mathcal{A}\to[0,R_{\max}]$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $d_{0}$ is the initial state distribution. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ induces a distribution over random trajectories, generated as $s_{0}\sim d_{0}$, $a_{t}\sim\pi(\cdot|s_{t})$, $r_{t}=R(s_{t},a_{t})$, $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. We use $\operatorname{Pr}_{\pi}[\cdot]$ and $\mathbb{E}_{\pi}[\cdot]$ to denote such a distribution and the expectation thereof. The performance of a policy is defined as $J(\pi):=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, which is in the range of $0,V_{\max}]$ where $V_{\max}:=R_{\max}/(1-\gamma)$.
@@
  • Model-based: Each OPE instance outputs an MDP $M_{i}$ and uses $J_{M_{i}}(\pi)$ as an estimate of $J(\pi)$. W.l.o.g. we assume $M_{i}$ only differs from $M^{\star}$ in the transition $P_{i}$, as handling different reward functions is straightforward. The task is to select $\hat{M}$ from $\mathcal{M}:=\{M_{i}\}_{i\in[m]}$, such that $J_{\hat{M}}(\pi)\approx J(\pi)$. Similar to the model-free case, we assume $M^{\star}\in\mathcal{M}$ in the derivations.
@@ 3.2 LSTD-Tournament
Besides the realizability of $Q^{\pi}$, the guarantee also depends on the invertibility of $A$, which can be viewed as a coverage condition, since $A$ changes with the data distribution $\mu$ [2, 3, 23]. In fact, in the on-policy setting, $\sigma_{\min}(A)$ can be shown to be lower-bounded away from 0 [42]; see Appendix C.5.
@@
The subscript “i,ji,j” makes it clear that the feature is created based on $Q_{i}$ and $Q_{j}$ as candidates, and we will use similar conventions for all quantities induced from $\phi_{i,j}$, e.g., $A_{i,j},b_{i,j}$, etc. Obviously, $Q^{\pi}$ is linear in $\phi_{i,j}$ with $\theta^{\star}\in\{[1,0]^{\top},[0,1]^{\top}\}$. Therefore, to choose between $Q_{i}$ and $Q_{j}$, we can calculate the LSTDQ loss of $1,0]^{\top}$ and $0,1]^{\top}$ under feature $\phi_{i,j}$ and choose the one with smaller loss. For $\theta=[1,0]^{\top}$, we have $A_{i,j}\theta-b_{i,j}=$
@@ 4 New Model-Based Selectors
    Naïve Baseline. A natural method is to use the model-prediction loss: any model $M$ is scored by $\mathbb{E}_{(s,a,s^{\prime})\sim\mu,\tilde{s}\sim P(\cdot|s,a)}[\|s^{\prime}-\tilde{s}\|$], where $P$ is the dynamics of $M$, and $\|\cdot\|$ is some norm (e.g., $\ell_{2}$ norm) that measures the distance between states. Despite its wide use in the model-based RL literature, the loss has a crucial caveat that it exhibits a bias towards more deterministic systems, which is problematic when the groundtruth environment is stochastic. We will see empirical evidence of this in Section 6 (see “mb_naive” in Figures 3 and 4).
@@ 5.1 The Protocol
    Candidate model list $\mathcal{M}=\{M_{i}\}_{i\in[m]}$.
@@
Our innovation here is to note that, for all the model-free algorithms we are interested in evaluating, they all access $Q_{M}^{\pi}$ exclusively through the value of $Q_{M}^{\pi}(s,a)$ and $Q_{M}^{\pi}(s^{\prime},\pi)$ for $(s,a,r,s^{\prime})$ in the offline dataset $\mathcal{D}$. That is, given $n$ data points in $\mathcal{D}$, we only need to know $2n$ scalar values about $Q_{M}^{\pi}$. Therefore, we propose to directly compute these values without explicitly representing $Q_{M}^{\pi}$, and each value can be easily estimated by averaging over multiple Monte-Carlo rollouts [50], i.e., $Q_{M}^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s,a_{0}=a]$ can be approximated by rolling out multiple trajectories starting from $(s,a)$ and taking actions according to $\pi$. For the model-based estimators proposed in Section 4, we need access to quantities in the form of $(\mathcal{T}_{M_{j}}^{\pi}Q_{M_{i}}^{\pi})(s,a)$ (see Appendix B). This value can also be obtained by Monte-Carlo simulation: (1) start in $(s,a)$ and simulate one step in $M_{j}$, then (2) switch to $M_{i}$, simulate from step 2 onwards and rollout the rest of the trajectory.
@@ 6 Exemplification of the Protocol
In this section we instantiate our protocol in the Gym Hopper environment to demonstrate its utility, while also providing preliminary empirical results for our algorithms. Our code is available at https://github.com/Coder-PAI/2025_neurips_model_selection_rl.git.
@@ 6.1 Experiment Setup and Main Results
Our experiments will be based on the Hopper-v4 environment [9]. To create a variety of environments, we add different levels of stochastic noise in the transitions and change the gravity constant (see Appendix D.1). Each environment is then parameterized by the gravity constant $\mathbf{g}$ and noise level $\mathbf{n}$. We consider arrays of such environments as the set of candidate simulator $\mathcal{M}$: in most of our results, we consider a “gravity grid” (denoted using MF.G in the figures) $\mathcal{M}_{\mathbf{g}}:=\{M_{\mathbf{g}}^{0}\ldots,M_{\mathbf{g}}^{14}\}$ (fixed noise level, varying gravity constants from 51-51 to 9-9) and a “noise grid” (MF.N) $\mathcal{M}_{\mathbf{n}}:=\{M_{\mathbf{n}}^{0}\ldots,M_{\mathbf{n}}^{14}\}$ (fixed gravity constant, varying noise level from $10$ to $100$). Each array contains 15 environments, though some subsequent results may only involve a subset of them (Section 6.2). Some of these simulators will also be treated as groundtruth environment $M^{\star}$, which determines the groundtruth performance of target policies and produces the offline dataset $\mathcal{D}$.
@@ 6.3 Data Coverage
In the previous subsection, we have seen how multiple experiment units that only differ in $\mathcal{M}$ can provide useful insights. Here we show that we can also probe the methods’ sensitivity to data coverage by looking at experiment units that only differ in the dataset $\mathcal{D}$. In Figure 5R, we take a previous experiment setting ($\mathcal{M}_{\mathbf{g}}$) and isolate a particular target policy $\pi$; then, we create two datasets: (1) $\mathcal{D}_{\pi}$ sampled using $\pi$; (2) $\mathcal{D}_{\textrm{off}}$ sampled using a policy that is created to be very different from the target policies and offer very little coverage (see Appendix D.2). Then, we run the algorithm with $\lambda$ fraction of data from $\mathcal{D}_{\pi}$ combined with $(1-\lambda)$ from $\mathcal{D}_{\textrm{off}}$; as predicted by theory, most methods perform better with better coverage (large $\lambda$), and performance degrades as $\lambda$ goes to 0.
@@ 7 Limitations, Future Directions, and Conclusion
As another potential future direction, LSTDQ can be viewed as an application of instrumental-variable regression in value-function estimation [8, 11], where the left $\phi$ in the moments of Eq.(3) plays the role of an instrument. While such a choice is standard and natural, it will be interesting to explore alternative choices of instruments and examine if they can lead to improved theoretical guarantees and practical performance in the model-selection problem.
Acknowledgements
Nan Jiang acknowledges funding support from NSF CNS-2112471, NSF CAREER IIS-2141781, Google Scholar Award, and Sloan Fellowship.
@@ Appendix A Other Related Works
Our theoretical guarantees state that the selected candidate function/model provides a $J(\pi)$ estimate that is close to the groundtruth, as if the algorithm knew which candidate is correct. This can be viewed as a form of adaptivity guarantees, which is the goal of several theoretical works on model selection. Su et al. [51] and Udagawa et al. [56] study adaptive model selection for OPE, but their approaches crucially rely on the importance sampling estimator, which we do not consider due to the focus on long-horizon tasks. Lee et al. [33] (who build on and improve upon the earlier works of Farahmand and Szepesvári [15] and Jiang et al. [24]) study model-selection of value functions, with a focus on the double-sampling issue and its relationship with Bellman completeness; these are also the very core consideration in our theoretical derivations. However, their approach treats the OPE instances (such as FQE) in a less black-box manner, and the goal is to select the “right” function class for the OPE algorithm (instead of directly selecting a final output function) from a nested series of classes, one of which is assumed to be Bellman-complete [6, 10] and have low generalization errors. [^4] Given the relationship between Bellman-completeness and bisimulation abstractions [10, Proposition 9], their setting is a natural generalization of Jiang et al. [24]’s setting of selecting a bisimulation from a nested series of state abstractions. This makes their approach less widely applicable than ours, and we cannot empirically evaluate their algorithm in our protocol as a consequence. Nevertheless, their results provide an interesting alternative and valuable insights to the model-selection problem, which are “complementary” to the BVFT line of work [62, 64] that we further develop. Similarly, Miyaguchi [41] considers the selection of Bellman operators, which can be instantiated as regression using different function classes, making their approach similar to Lee et al. [33] in spirit.
@@
Experimental Protocol
The closest to our experimental protocol is the work of Sajed et al. [50] who also uses Monte-Carlo rollouts to produce unbiased estimates of value functions on individual states. Other than this point, their work is largely orthogonal to ours as they focus on establishing high-confidence bounds for the estimated value, assuming the simulator is the groundtruth environment, whereas we draw Monte-Carlo trajectories from different simulators to facilitate the model-selection problem.
@@ Appendix B Details of Model-based Selectors
[…]
@@
Despite its wide use and simplicity [17, 45, 18], the method has major caveats: first, the distance metric $d(\cdot,\cdot)$ is a design choice. When the state is represented as a real-valued vector, it is natural to use the $\ell_{2}$ distance as $d(\cdot,\cdot)$, which changes if we simply normalize/rescale the coordinates. Second, the metric is biased for stochastic environments as discussed in prior works [21, 59], which we also demonstrate in the experiment section (Section 6); essentially this is a version of the double-sampling issue but for the model-based setting [5]. As a minimal counterexample, suppose $\mathcal{S}\subset\mathbb{R}^{2}$ and we focus on the transition distribution from a particular $(s,a)$ pair where groundtruth is uniform over 4 points $(\pm 1,\pm 1)$. The expected loss of the groundtruth model itself, as in Eq.(12), is $1+\sqrt{2}/2$; this is higher than a wrong model that always predicts 0 deterministically, which yields a loss of $\sqrt{2}$.